OpenStudy (anonymous):

The tangent line to the graph of f(x) = x^2 at the point (x, y) has slope = 2x. The equation of the tangent line to the graph of f(x) = x^2 at x = -2 is y =?

OpenStudy (anonymous):

y=-4x

OpenStudy (anonymous):

slope is -4

OpenStudy (anonymous):

the slope is -4 right ,but y=-4x?

OpenStudy (anonymous):

because f(-2) = -4 right?

OpenStudy (anonymous):

is that correct 14yamaka?

OpenStudy (anonymous):

can you show me how you got that, yamaka?

OpenStudy (anonymous):

sry, just got back XD ok

OpenStudy (anonymous):

I meant slope is four. This is how we get to the answer: FIrst, plug x=-2 in as that is the point where the line and the f(x)=x^2 cross. f(-2)=(x)^2 f(-2)=4 (-2, 4) Now, we also know that slope is 2x. THerefore, slope is -4. So, y=-4x+b 4=-4(-2)+b b=-4 y=-4x-4